### v2

@@ 5.5 GST Circuit Design
However, the uncertainty of the extracted Pauli $X$ and Pauli $Y$ errors from GST is very large due to the shallow depth of the GST circuits (128 at maximum), which hinders the exact characterization of the noise-bias. Therefore, we employ the DRB protocol with a maximum circuit depth of 2000 to explore the noise structure of KCQ and demonstrate that the performance of detuned-KCQ crosses the fault-tolerant threshold of the XZZX surface code[19, 42].
5.6 $\mathbb{D}_{8}$ Dihedral Group Randomized Benchmarking Design
The bit-flip error and phase-flip error, as well as the noise bias are extracted by the DRB protocol introduced in references[46, 42] with a scaling factor to compensate the effect of noise-free $X(\pi)$ gate. We will describe the protocol and illustrate the extraction of the scaling factor by numerical simulations.
@@ 5.6 $\mathbb{D}_{8}$ Dihedral Group Randomized Benchmarking Design
Figure 12: DRB survival probability fitting. Circuits and fitting curve examples of the exponential decay in the $\mathbb{D}_{8}$ dihedral randomized benchmarking protocol when a, c preparing and measuring the states in $Z$ axis, and b, d preparing and measuring the states in $X$ axis.
The scaling factor between the errors of $Z(\theta)$ rotations and the extracted errors from DRB protocol is derived by numerical simulation with qiskit[74] package. We introduce various phase-flip errors $p_{\text{pf, \text{real}}}$ ranging from 00 to $3\%$ and bit-flip errors $p_{\text{bf, real}}=0.02p_{\text{pf, real}}$ to $Z(\theta)$ rotations, while keeping the $X(\pi)$ gate noise-free, and then simulate the DRB experiments to extract the bit-flip $p_{\text{pf, extracted}}$ and phase-flip errors $p_{\text{pf, extracted}}$. As shown in Fig. 13, the extracted errors are proportional to the errors on the $Z(\theta)$ rotations. By fitting the simulation results with a linear function, we extracted the scaling factor between them, where $p_{\text{pf, real}}=1.02p_{\text{pf, extracted}}$ and $p_{\text{bf, real}}=1.07p_{\text{bf, extracted}}$.
@@
5.7 Description and Analysis of the Error Channels
We model the error channels of KCQ as a combination of coherent Pauli rotation errors and incoherent stochastic Pauli errors described by a Pauli transfer matrix $\mathcal{E}=e^{\mathbf{L}}$ with the generator $\mathbf{L}=h_{x}\mathbf{H_{x}}+h_{y}\mathbf{H_{y}}+h_{z}\mathbf{H_{z}}+p_{x}\mathbf{P_{x}}+p_{y}\mathbf{P_{y}}+p_{z}\mathbf{P_{z}}$. The coherent components in the error generator are defined in Eq. 16, and the incoherent components are defined in Eq. 17.[75, 76]
@@ 5.7 Description and Analysis of the Error Channels
The coherent error manifests as the off-diagonal elements of the Pauli transfer matrix, which can be effectively removed by the Pauli twirling process. The Pauli twirling process randomly insert Pauli gates $P\in\mathbb{P}$ while compiling circuits, and all the Pauli gates can be executed with error-free virtual $X$ gate, high-fidelity bias-preserving $Z$ gate or their combination. Therefore, Pauli twirled error channels introduces minimal extra noise with preserved the noise-bias property, and are reported to be statistically the same as the full error channels in $\mathtt{XZZX}$ surface code[19].

### v3

@@ 5 GST Circuit Design
However, the uncertainty of the extracted Pauli $X$ and Pauli $Y$ errors from GST is very large due to the shallow depth of the GST circuits (128 at maximum), which hinders the exact characterization of the noise-bias. Therefore, we employ the DRB protocol with a maximum circuit depth of 2000 to explore the noise structure of KCQ and demonstrate that the performance of detuned-KCQ crosses the fault-tolerant threshold of the XZZX surface code(22, 41).
6 $\mathbb{D}_{8}$ Dihedral Group Randomized Benchmarking Design
The bit-flip error and phase-flip error, as well as the noise bias are extracted by the DRB protocol introduced in references(40, 41) with a scaling factor to compensate the effect of noise-free $X(\pi)$ gate. We will describe the protocol and illustrate the extraction of the scaling factor by numerical simulations.
@@ 6 $\mathbb{D}_{8}$ Dihedral Group Randomized Benchmarking Design
Figure 12: DRB survival probability fitting. Circuits and fitting curve examples of the exponential decay in the $\mathbb{D}_{8}$ dihedral randomized benchmarking protocol when a, c preparing and measuring the states in $Z$ axis, and b, d preparing and measuring the states in $X$ axis. The error bars indicate the shot noise from 1024 repeated measurements.
The scaling factor between the errors of $Z(\theta)$ rotations and the extracted errors from DRB protocol is derived by numerical simulation with qiskit(76) package. We introduce various phase-flip errors $p_{\text{pf, \text{real}}}$ ranging from 0 to $3\%$ and bit-flip errors $p_{\text{bf, real}}=0.02p_{\text{pf, real}}$ to $Z(\theta)$ rotations, while keeping the $X(\pi)$ gate noise-free, and then simulate the DRB experiments to extract the bit-flip $p_{\text{pf, extracted}}$ and phase-flip errors $p_{\text{pf, extracted}}$. As shown in Fig. 13, the extracted errors are proportional to the errors on the $Z(\theta)$ rotations. By fitting the simulation results with a linear function, we extracted the scaling factor between them, where $p_{\text{pf, real}}=1.02p_{\text{pf, extracted}}$ and $p_{\text{bf, real}}=1.07p_{\text{bf, extracted}}$.
@@
7 Description and Analysis of the Error Channels
We model the error channels of KCQ as a combination of coherent Pauli rotation errors and incoherent stochastic Pauli errors described by a Pauli transfer matrix $\mathcal{E}=e^{\mathbf{L}}$ with the generator $\mathbf{L}=h_{x}\mathbf{H_{x}}+h_{y}\mathbf{H_{y}}+h_{z}\mathbf{H_{z}}+p_{x}\mathbf{P_{x}}+p_{y}\mathbf{P_{y}}+p_{z}\mathbf{P_{z}}$. The coherent components in the error generator are defined in Eq. 16, and the incoherent components are defined in Eq. 17.(77, 78)
@@ 7 Description and Analysis of the Error Channels
The coherent error manifests as the off-diagonal elements of the Pauli transfer matrix, which can be effectively removed by the Pauli twirling process. The Pauli twirling process randomly insert Pauli gates $P\in\mathbb{P}$ while compiling circuits, and all the Pauli gates can be executed with error-free virtual $X$ gate, high-fidelity bias-preserving $Z$ gate or their combination. Therefore, Pauli twirled error channels introduces minimal extra noise with preserved the noise-bias property, and are reported to be statistically the same as the full error channels in $\mathtt{XZZX}$ surface code(22).